# Continual Learning for Encoder-only Language Models via a Discrete Key-Value Bottleneck

### v1

@@ 9 Limitations
Our study focused solely on encoder-only language models. While this raises questions whether our results could generalize to other model architectures, our choice was motivated by their preference for supervised fine-tuning scenarios where the balance between performance and computational efficiency is crucial. Our experiments were also limited to fine-tuning for text classification-based downstream tasks. Consequently it remains uncertain whether our results extend to other NLP tasks, such as semantic search, entity extraction or machine translation.
Appendix A Appendix
A.1 Extended Related Work
@@
This family is unique to the field of NLP. These methods are based on task-specific instructions given to encoder-decoder or decoder only language models when a new task is encountered. While some methods in this family show promising knowledge transfer capabilities (Scialom et al., 2022; Yin et al., 2022), without explicit fine-tuning they are limited by the knowledge acquired in the pre-training phase.
A.2 Extended Experimental Setup
@@
For all model backbones in our experiments, we use the BERT-base model from Huggingface [^2] https://huggingface.co/bert-base-uncased and use cross-entropy loss as our objective function. We base our discrete-key-value bottleneck implementation on the vector-quantize-pytorch [^3] https://github.com/lucidrains/vector-quantize-pytorch package. In the pre-experiments, we truncate each input sample to 256 tokens. For the main continual learning experiments, we rely on the PyContinual [^4] https://github.com/ZixuanKe/PyContinual framework and reuse their implementations and hyperparameters on the baseline methods. To remain comparable to other studies using the PyContinual framework, we kept the default preprocessing steps, used a maximum token length of 128, and applied the default convolutional decoder of the baseline models. For the single-head class incremental learning experiments we use a fixed randomized sequence order when creating the increments, and used a token length of 256. The source code for our experiments alongside the models can be found at anonymized GitHub link.
@@
As part of our pre-experiments, we also conducted a hyperparameter search and a sensitivity analysis on the bottleneck parameters. Outside of the selected hyperparameters and bottleneck parameters, all other configurations remained fixed during the search. Our experiments use the BERT-base architecture with a hidden size of 768. For the optimizer, we chose AdamW with a weight decay of 0.01. The dropout rate for the parametric decoder was set to 0.1. For the reference fully fine-tuned BERT numbers we reused the hyperparameters reported in Galke and Scherp (2022), for the frozen BERT variant we relied on the parametric DKVB variant hyperparameters with mean pooling. During fine-tuning, we carefully optimized the models on both datasets using grid-search-based manual tuning. A search space for the selected hyperparameters was defined, specifically we chose the batch size from $\{8,16,32\}$, the number of epochs from $\{5,10\}$, the learning rate for the values layer from $\{1e-1,1e-2,1e-3\}$, and the decoder learning rate from $\{1e-3,1e-4,1e-5\}$. The best performing (based on the validation loss) configurations for each architecture variant can be seen in Table 4. The hyperparameters were reused for the continual learning main experiments.
In the single-head class incremental learning experiments we conducted an additional manual hyperparameter search for the DKVB variants and found a batch size of 16 with a global learning rate of $21e-$ 2 to be the best performing. For EWC we set the lambda parameter to 5000. In the DER++ model we used a memory buffer size of 256 and set the sampling reate in each increment to 16.
@@
DKVB-P hidden Before CLS 5 32 $21e-$ 2 $41e-$ 4 10 16 $11e-$ 1 $31e-$ 3
DKVB-P hidden Before Mean 5 32 $11e-$ 1 $41e-$ 4 10 16 $21e-$ 2 $41e-$ 4
DKVB-P hidden After CLS 10 16 $21e-$ 2 $41e-$ 4 5 16 $21e-$ 2 $41e-$ 4
DKVB-P hidden After Mean 10 16 $21e-$ 2 $31e-$ 3 5 16 $21e-$ 2 $31e-$ 3
DKVB-P token After CLS 10 16 $21e-$ 2 $31e-$ 3 10 16 $21e-$ 2 $31e-$ 3
DKVB-P token After Mean 10 16 $21e-$ 2 $31e-$ 3 10 16 $21e-$ 2 $31e-$ 3
DKVB-NP hidden Before CLS 5 32 $11e-$ 1 - 5 32 $11e-$ 1 -
DKVB-NP hidden Before Mean 5 32 $11e-$ 1 - 10 32 $11e-$ 1 -
DKVB-NP hidden After CLS 10 16 $11e-$ 1 - 5 16 $11e-$ 1 -
DKVB-NP hidden After Mean 10 32 $11e-$ 1 - 10 16 $11e-$ 1 -
DKVB-NP token After CLS 10 32 $11e-$ 1 - 10 32 $21e-$ 2 -
DKVB-NP token After Mean 10 32 $11e-$ 1 - 10 32 $21e-$ 2 -
A.3 Bottleneck Parameters
@@ A.4 Extended Results
Results obtained using the RoBERTa (Table 5) and DistilBERT (Table 6) models demonstrate comparable performance patterns to those observed with BERT. Interestingly, DistilBERT produced slightly better accuracy on most architecture variants compared to the other two model. However the highest performance was consistently seen with mean pooling after the bottleneck on all three models. DistilBERT’s improved performance can likely be attributed to differences in its tokenization and pooling implementation compared to other models.
@@
Dataset: R8 Dataset: 20ng Decoder Segmentation Pooling Before Pooling After Pooling Before Pooling After CLS Mean CLS Mean CLS Mean CLS Mean Parametric hidden $49.48_{\textit{0.05}}$ $91.36_{\textit{0.40}}$ $91.73_{\textit{0.81}}$ $\mathbf{94.25_{\textit{0.26}}}$ $51.05_{\textit{0.43}}$ $56.63_{\textit{0.39}}$ $52.18_{\textit{1.11}}$ $\mathbf{75.08_{\textit{0.21}}}$ token - - $90.02_{\textit{1.05}}$ $94.05_{\textit{0.31}}$ $19.86_{\textit{1.03}}$ $27.30_{\textit{0.92}}$ Non Parametric hidden $49.45_{\textit{0.02}}$ $92.05_{\textit{0.29}}$ $74.53_{\textit{1.74}}$ $93.04_{\textit{0.20}}$ $56.83_{\textit{0.35}}$ $60.42_{\textit{0.35}}$ $53.10_{\textit{1.37}}$ $70.33_{\textit{0.78}}$ token - - $58.74_{\textit{0.92}}$ $66.33_{\textit{0.74}}$ $9.89_{\textit{0.66}}$ $12.51_{\textit{1.67}}$ RoBERTa (frozen) w/o DKVB $94.29_{\textit{0.17}}$ $69.42_{\textit{0.30}}$ RoBERTa w/o DKVB $97.54_{\textit{0.51}}$ $83.36_{\textit{0.30}}$
@@
Dataset: R8 Dataset: 20ng Decoder Segmentation Pooling Before Pooling After Pooling Before Pooling After CLS Mean CLS Mean CLS Mean CLS Mean Parametric hidden $90.22_{\textit{0.30}}$ $92.17_{\textit{0.25}}$ $90.46_{\textit{0.44}}$ $\mathbf{96.09_{\textit{0.21}}}$ $56.26_{\textit{0.53}}$ $60.24_{\textit{0.20}}$ $60.26_{\textit{0.31}}$ $\mathbf{79.79_{\textit{0.49}}}$ token - - $89.24_{\textit{0.93}}$ $94.78_{\textit{0.86}}$ $45.45_{\textit{1.13}}$ $68.06_{\textit{0.90}}$ Non Parametric hidden $89.79_{\textit{0.24}}$ $92.02_{\textit{0.38}}$ $90.92_{\textit{0.22}}$ $95.09_{\textit{0.27}}$ $60.73_{\textit{0.48}}$ $59.98_{\textit{0.30}}$ $61.43_{\textit{0.50}}$ $75.11_{\textit{0.44}}$ token - - $66.37_{\textit{0.49}}$ $72.65_{\textit{0.24}}$ $12.04_{\textit{0.51}}$ $18.70_{\textit{1.03}}$ DistilBERT (frozen) w/o DKVB $94.62_{\textit{0.16}}$ $68.56_{\textit{0.38}}$ DistilBERT w/o DKVB $97.83_{\textit{0.24}}$ $83.84_{\textit{0.23}}$
@@
where $R\in\mathbb{R}^{T\times T}$ is the results matrix of an incremental learning scenario with $T$ tasks, where each entry $R_{i,j}$ being the test accuracy on task $j$ after training on task $i$ Lopez-Paz and Ranzato (2017).We report the BWT numbers on the three continual learning scenarios in Table 7..
@@
CL Method Model DIL (DSC) CIL (20NG) TIL (4GLUE)
NCL BERT $0.29$ $-29.77$ $-20.00$
NCL BERT (frozen) $-0.10$ $-0.38$ $-6.70$
NCL Adapter-BERT $0.39$ $-20.01$ $-16.05$
DER++ BERT (frozen) $-0.58$ $-27.11$ $-7.05$
EWC BERT (frozen) $0.06$0.27 $-10.84$
OWM BERT (frozen) $0.26$ $-15.44$ $-8.23$
CTR Adapter-BERT 0.49 $-0.50$6.19
DKVB-NP Incr. BERT (frozen) $-3.06$ $-27.73$ $-21.30$
@@
DKVB-NP Generic BERT (frozen) $-1.17$ $-0.29$ $-7.97$
DKVB-P Incr. BERT (frozen) $-4.88$ $-29.05$ $-20.99$
DKVB-P Oracle BERT (frozen) $-1.02$ $-0.41$ $-20.56$
DKVB-P Generic BERT (frozen) $-6.24$ $-4.94$ $-16.00$
Single-head Class Incremental Learning
In Table 8 and 9 we report the extended results of the single-head class incremental learning experiments. The number of test samples column refers to the number of samples in the test set that possess the same class/classes as the training data in a given increment.
@@
Increment # Test Samples BERT BERT-frozen BERT-frozen DER++ BERT-frozen EWC DKVB-NP Incr. DKVB-NP Oracle DKVB-NP Wiki
@@
Increment # Test Samples BERT BERT-frozen BERT-frozen DER++ BERT-frozen EWC DKVB-NP Incr. DKVB-NP Oracle DKVB-NP Wiki

### v2

@@ 9 Limitations
Our study focuses solely on encoder-only language models. While this raises questions about whether our results could generalize to other model architectures, our choice was motivated by their preference for supervised fine-tuning scenarios where the balance between performance and computational efficiency is crucial. Our experiments were also limited to fine-tuning for classification-based downstream tasks. Consequently, it remains to be investigated whether our results extend to other NLP tasks, such as semantic search, entity extraction, or machine translation.
Supplementary Materials
Appendix A Extended Related Work
@@
This family is unique to the field of NLP. These methods are based on task-specific instructions given to encoder-decoder or decoder only language models when a new task is encountered. While some methods in this family show promising knowledge transfer capabilities (Scialom et al., 2022; Yin et al., 2022; Razdaibiedina et al., 2023), without explicit fine-tuning they are mostly limited by the knowledge acquired in the pre-training phase.
Appendix B Extended Experimental Setup
@@
For all model backbones in our experiments, we use the BERT-base model from Huggingface [^3] https://huggingface.co/bert-base-uncased and use cross-entropy loss as our objective function. We base our discrete-key-value bottleneck implementation on the vector-quantize-pytorch [^4] https://github.com/lucidrains/vector-quantize-pytorch package. In the pre-experiments, we truncate each input sample to 256 tokens. For the main continual learning experiments, we rely on the PyContinual [^5] https://github.com/ZixuanKe/PyContinual framework and reuse their implementations and hyperparameters on the baseline methods. To remain comparable to other studies using the PyContinual framework, we kept the default preprocessing steps, used a maximum token length of 128, and applied the default convolutional decoder of the baseline models. For the single-head class incremental learning experiments we use a fixed randomized sequence order when creating the increments, and used a token length of 256. The source code for our experiments alongside the models can be found at github.com/drndr/dkvb_nlp.
@@
[…]
In the single-head class incremental learning experiments we conducted an additional manual hyperparameter search for the DKVB variants and found a batch size of 16 with a global learning rate of $1e-2$ to be the best performing. For EWC we set the lambda parameter to 5,000. In the DER++ model we used a memory buffer size of 256 and set the sampling reate in each increment to 16.
@@
DKVB-P hidden Before CLS 5 32 $1e\text{-}2$ $1e\text{-}4$ 10 16 $1e\text{-}1$ $1e\text{-}3$
DKVB-P hidden Before Mean 5 32 $1e\text{-}1$ $1e\text{-}4$ 10 16 $1e\text{-}2$ $1e\text{-}4$
DKVB-P hidden After CLS 10 16 $1e\text{-}2$ $1e\text{-}4$ 5 16 $1e\text{-}2$ $1e\text{-}4$
DKVB-P hidden After Mean 10 16 $1e\text{-}2$ $1e\text{-}3$ 5 16 $1e\text{-}2$ $1e\text{-}3$
DKVB-P token After CLS 10 16 $1e\text{-}2$ $1e\text{-}3$ 10 16 $1e\text{-}2$ $1e\text{-}3$
DKVB-P token After Mean 10 16 $1e\text{-}2$ $1e\text{-}3$ 10 16 $1e\text{-}2$ $1e\text{-}3$
DKVB-NP hidden Before CLS 5 32 $1e\text{-}1$ 5 32 $1e\text{-}1$
DKVB-NP hidden Before Mean 5 32 $1e\text{-}1$ 10 32 $1e\text{-}1$
DKVB-NP hidden After CLS 10 16 $1e\text{-}1$ 5 16 $1e\text{-}1$
DKVB-NP hidden After Mean 10 32 $1e\text{-}1$ 10 16 $1e\text{-}1$
DKVB-NP token After CLS 10 32 $1e\text{-}1$ 10 32 $1e\text{-}2$
DKVB-NP token After Mean 10 32 $1e\text{-}1$ 10 32 $1e\text{-}2$
B.1 Bottleneck Parameters
@@ Appendix C Extended Results
Results obtained using the RoBERTa (Table 6) and DistilBERT (Table 7) models demonstrate comparable performance patterns to those observed with BERT. Interestingly, DistilBERT produced slightly better accuracy on most architecture variants compared to the other two models. However the highest performance was consistently seen with mean pooling after the bottleneck on all three models. DistilBERT’s improved performance can likely be attributed to differences in its tokenization and pooling implementation compared to other models.
@@
Dataset: R8 Dataset: 20ng
Decoder Segmentation Pooling Before Pooling After Pooling Before Pooling After
CLS Mean CLS Mean CLS Mean CLS Mean
Parametric hidden $49.48_{\textit{0.05}}$ $91.36_{\textit{0.40}}$ $91.73_{\textit{0.81}}$ $\mathbf{94.25_{\textit{0.26}}}$ $51.05_{\textit{0.43}}$ $56.63_{\textit{0.39}}$ $52.18_{\textit{1.11}}$ $\mathbf{75.08_{\textit{0.21}}}$
token - - $90.02_{\textit{1.05}}$ $94.05_{\textit{0.31}}$ $19.86_{\textit{1.03}}$ $27.30_{\textit{0.92}}$
Non Parametric hidden $49.45_{\textit{0.02}}$ $92.05_{\textit{0.29}}$ $74.53_{\textit{1.74}}$ $93.04_{\textit{0.20}}$ $56.83_{\textit{0.35}}$ $60.42_{\textit{0.35}}$ $53.10_{\textit{1.37}}$ $70.33_{\textit{0.78}}$
token - - $58.74_{\textit{0.92}}$ $66.33_{\textit{0.74}}$ $9.89_{\textit{0.66}}$ $12.51_{\textit{1.67}}$
RoBERTa (frozen) w/o DKVB $94.29_{\textit{0.17}}$ $69.42_{\textit{0.30}}$
RoBERTa w/o DKVB $97.54_{\textit{0.51}}$ $83.36_{\textit{0.30}}$
@@
Dataset: R8 Dataset: 20ng
Decoder Segmentation Pooling Before Pooling After Pooling Before Pooling After
CLS Mean CLS Mean CLS Mean CLS Mean
Parametric hidden $90.22_{\textit{0.30}}$ $92.17_{\textit{0.25}}$ $90.46_{\textit{0.44}}$ $\mathbf{96.09_{\textit{0.21}}}$ $56.26_{\textit{0.53}}$ $60.24_{\textit{0.20}}$ $60.26_{\textit{0.31}}$ $\mathbf{79.79_{\textit{0.49}}}$
token - - $89.24_{\textit{0.93}}$ $94.78_{\textit{0.86}}$ $45.45_{\textit{1.13}}$ $68.06_{\textit{0.90}}$
Non Parametric hidden $89.79_{\textit{0.24}}$ $92.02_{\textit{0.38}}$ $90.92_{\textit{0.22}}$ $95.09_{\textit{0.27}}$ $60.73_{\textit{0.48}}$ $59.98_{\textit{0.30}}$ $61.43_{\textit{0.50}}$ $75.11_{\textit{0.44}}$
token - - $66.37_{\textit{0.49}}$ $72.65_{\textit{0.24}}$ $12.04_{\textit{0.51}}$ $18.70_{\textit{1.03}}$
DistilBERT (frozen) w/o DKVB $94.62_{\textit{0.16}}$ $68.56_{\textit{0.38}}$
DistilBERT w/o DKVB $97.83_{\textit{0.24}}$ $83.84_{\textit{0.23}}$
@@
where $R\in\mathbb{R}^{T\times T}$ is the results matrix of an incremental learning scenario with $T$ tasks, where each entry $R_{i,j}$ being the test accuracy on task $j$ after training on task $i$ Lopez-Paz and Ranzato (2017).We report the BWT numbers on the three continual learning scenarios in Table 8.
@@
CL Method Model DIL CIL TIL
(DSC) (20NG) (4GLUE)
NCL BERT 0.290.290.29 29.77-29.77- 29.77 20.00-20.00- 20.00
NCL BERT (frozen) 0.10-0.10- 0.10 0.38-0.38- 0.38 6.70-6.70- 6.70
NCL Adapter-BERT 0.390.390.39 20.01-20.01- 20.01 16.05-16.05- 16.05
DER++ BERT (frozen) 0.58-0.58- 0.58 27.11-27.11- 27.11 7.05-7.05- 7.05
EWC BERT (frozen) 0.060.060.06 -0.27 10.84-10.84- 10.84
OWM BERT (frozen) 0.260.260.26 15.44-15.44- 15.44 8.23-8.23- 8.23
CTR Adapter-BERT 0.49 0.50-0.50- 0.50 -6.19
DKVB-NP Incremental BERT (frozen) 3.06-3.06- 3.06 27.73-27.73- 27.73 21.30-21.30- 21.30
@@
DKVB-NP Generic BERT (frozen) 1.17-1.17- 1.17 0.29-0.29- 0.29 7.97-7.97- 7.97
DKVB-P Incremental BERT (frozen) 4.88-4.88- 4.88 29.05-29.05- 29.05 20.99-20.99- 20.99
DKVB-P Oracle BERT (frozen) 1.02-1.02- 1.02 0.41-0.41- 0.41 20.56-20.56- 20.56
DKVB-P Generic BERT (frozen) 6.24-6.24- 6.24 4.94-4.94- 4.94 16.00-16.00- 16.00
@@
Increment # Test Samples BERT BERT-frozen BERT-frozen DER++ BERT-frozen EWC DKVB-NP Incremental DKVB-NP Oracle DKVB-NP Wiki
@@
Increment # Test Samples BERT BERT-frozen BERT-frozen DER++ BERT-frozen EWC DKVB-NP Incremental DKVB-NP Oracle DKVB-NP Wiki